OpenStudy (anonymous):

Please can you help me sin^2A + sin^2B + sin^2C = 2 + 2cosAcosBcosC

Miracrown (miracrown):

What are the directions to this problem? @sladi4koto_rumi

OpenStudy (anonymous):

If A+B+C=180 prove that sin^2A + sin^2B + sin^2C = 2 + 2cosAcosBcosC

OpenStudy (anonymous):

ummm write sin^2A = (1-cos(2A))/2 and similar to other and try to solve?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/List_of_trigonometric_identities go down down down :P

OpenStudy (anonymous):

to power reduction formula

OpenStudy (anonymous):

(1-cos(2A))/2+(1-cos(2B))/2+(1-cos(2C))/2 solveee it?? atleast tell me where are u stuck o_o

OpenStudy (anonymous):

i think she is away? sigh

OpenStudy (anonymous):

(3- (cos2A + cos2B +cos2C) )/2 = (3 - ( -4cosAcosBcosC - 1))/2 = (4+4cosAcosBcosC)/2 = 2 + 2cosAcosBcosC right?

OpenStudy (anonymous):

tell me whats the formula for cos2A + cos2B

Miracrown (miracrown):

let me remind you some formulas which are going to be very useful cos(x+y) + cos(x-y) = 2 cos(x)cos(y) so let's start from the right side 2 + 2cos(A)cos(B)cos(C) = 2 + (cos(A+B) + cos(A-B))cos(C) and now from: A + B + C = 180, we can write that: A + B = 180 - C and we plug it 2 + (cos(180-C) + cos(A-B))cos(C) and cos(180-C) = -cos(C) and we use it 2 + (-cos(C) + cos(A-B))cos(C) now we distribute it 2 - cos^2(C) + cos(A-B)cos(C) we know use for the last C, C = 180 - A - B 2 - cos^2(C) + cos(A-B)cos(180 - A - B) and we apply again for the last term and as cos(180 - A - B) = - cos(A + B) we get 2 - cos^2(C) - cos(A-B)cos(A+B) now we apply again the first property 2cos(x)cos(y) = cos(x + y) + cos(x - y) in this case x = A - B and y = A + B cos(A - B)cos(A + B) = ( cos(A - B + A + B) + cos(A - B - A - B) ) /2 so let's plug it first let's reduce it cos(A - B)cos(A + B) = ( cos(2A) + cos(-2B) ) / 2 and as cos(- X ) = cos( x) we have cos(A - B)cos(A + B) = ( cos(2A) + cos(2B) ) / 2 2 - cos^2(C) - ( cos(2A) + cos(2B) ) /2 and here you just need to apply the definition of the double angle cos(2X) = cos^2(X) - sin^2(X) = 1 - 2sin^2(X) let's apply it 2 - cos^2(C) - ( 1 - 2sin^2(A) + 1 - 2sin^2(B) ) /2 and we apply the Pythagorean theore for the cos^2(C) cos^2(C) = 1 - sin^2(C) then we get: 2 - ( 1 - sin^2(C) ) - ( 2 - 2 sin^2(A) - 2sin^2(B) ) /2 distributing the signs 2 - 1 + sin^2(C) - 1 + sin^2(A) + sin^2(B) and as you can see 2 - 1 - 1 = 0 and we get the left side of the original expression: sin^2(A) + sin^2(B) + sin^2(C)

OpenStudy (anonymous):

O_O yes the hot chick solved the answer

OpenStudy (anonymous):

i mean the question

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

i mean proved it

Miracrown (miracrown):

No worries, good luck! @sladi4koto_rumi

OpenStudy (anonymous):

thank you

Miracrown (miracrown):

=)

ganeshie8 (ganeshie8):

Method 2 : starting from left side sin^2A + sin^2B + sin^2C 2 + sin^2C - cos^2A - cos^2B 2 + sin^2(A+B) - cos^2A - cos^2B ... 2 - 2cosAcosB[cosAcosB - sinAsinB] 2 - 2cosAcosB[cos(A+B)] 2 + 2cosAcosBcosC

ganeshie8 (ganeshie8):

Method 3 : take both sides try it :)